# HMT: Hierarchical Memory Transformer for Long Context Language Processing

### v2

@@ 5 Experiment
Table 3: Training configurations for the backbone models (OPT 350M, Mamba 370M, OPT 2.7B, RWKV 3B, OpenLlamaV2 3B, Llama 2 7B, Yi-6B-200K, and Mistral 7B) and the modified model after applying RMT and HMT. S1 and S2 indicate the first stage and the second stage of multi-stage training for HMT. 4 AMD MI210 GPUs cannot train larger models. The given learning rate is the starting learning rate, and will decay by factor of 0.7 for every 100 steps. The batch size is 2.
@@
Model Input Length Segment Length Learning Rate Training Steps
OPT 350M 2048 1024 1e-5 700
@@
We evaluate HMT and RMT with the same backbone models. Both HMT and RMT are benchmarked with a variety of backbone models including OPT 350M, OPT 2.7B (Zhang et al., 2022), OpenLlamaV2 (Geng & Liu, 2023) 3B, RWKV 3B (Peng et al., 2023), and Llama 2 7B (Touvron et al., 2023) models as the backbones. Moreover, we test several models with a long maximum context (Mamba 370M (Gu & Dao, 2023), Yi-6B-200K (Young et al., 2024), and Mistral 7B (Jiang et al., 2023)) to demonstrate the benefit HMT has on effectiveness and memory consumption. For general language modeling tasks, we select OPT 350M, OPT 2.7B, and OpenLlamaV2 3B as the representative of context-constrained models (models with a short maximum context window) and RWKV 3B and Mamba 370M for recurrent models. We evaluate the recurrent backbone model since we believe that HMT can be beneficial to these models that can already process infinitely long inputs. For the question-answering tasks, we employ Llama 2 7B as the backbone model. The largest model that our GPUs can handle has 7B parameters, but HMT produces consistently better performances over all these models to show that it can potentially be applied to larger models for an effectiveness boost. All models are trained and tested on 4 AMD MI210 GPUs. The baselines are evaluated using the sliding window attention for context-constrained models to control the memory consumption for a fair comparison. Due to the limited VRAM of GPUs, we shrink the segment length for the larger backbone model. Also, increasing memory token size does not improve the effectiveness as Bulatov et al. suggested. Thus, both RMT and HMT apply memory tokens with a length of 1. For RMT, we utilize the maximum BPTT unroll depth with the best effectiveness that the GPUs can handle. HMT is trained with the multi-stage training technique illustrated in Section 4.2. The first stage (S1) is trained with 2 segments, and the second stage (S2) is trained with the maximum BPTT unroll depth that the GPUs can manage. We select the learning rate for HMT, RMT, and the baseline to optimize the effectiveness. Furthermore, HMT preserves 32 tokens from the previous segment as the sensory memory. All models are trained with 700 steps in total, which is sufficient to converge the training loss. Table 3 recaps the training configurations of different models in our experiments.
For general language tasks, models are tested for next token generation tasks with Wikitext-103 (Merity et al., 2016) and PG-19 (Rae et al., 2019) datasets. Wikitext-103 contains general knowledge from Wikipedia where each term definition has 2 - 3k words. Processing inputs with multiple definitions can simulate context switches. PG-19 is a benchmark that includes a set of books published before 1919 from the Project Gutenberg library, with an average length of 69k tokens. Samples will be concatenated or split into chunks with specified input lengths to investigate the relationships between input length and the effectiveness of the model. For question-answering tasks, we chose PubMedQA (Jin et al., 2019), which is a biomedical question-answering dataset with corresponding contexts. We artifact the long-context dataset as the following: (1) select $M$ question-context-answer tuples from the dataset. Let this set of tuples be $\{(C_{0},Q_{0},A_{0}),(C_{1},Q_{1},A_{1}),\dots,(C_{T},Q_{T},A_{T})\}$, where $C_{n}$ are contexts, $Q_{n}$ are questions, $A_{n}$ are answers for $0\geq n\geq M$. Answers can be either long answers with detailed reasoning or short answers (either “yes”, “no”, or “maybe”). (2) Concatenate all the contexts from each tuple to form a long context and append the questions and answers for each tuple. This will create $M$ sequences: ${C_{0}C_{1}\dots C_{T}Q_{0}A_{0}}$, ${C_{0}C_{1}\dots C_{T}Q_{1}A_{1}},\dots$, ${C_{0}C_{1}\dots C_{T}Q_{T}A_{T}}$. By controlling the value of $M$, we can determine the fraction of useful information in the context for each question and better understand the filtering and selection ability of HMT and the baseline model.
Figure 4: Effectiveness of HMT, RMT, and three baseline models (OPT 2.7B, RWKV 3B, OpenLlamaV2 3B) with the Wikitext-103 dataset. HMT outperforms RMT by 13.0% for OPT and 10.8% for OpenLlamaV2. For RWKV, HMT can even boost the effectiveness by 16.5%, while RMT worsens the effectiveness.
Figure 5: Effectiveness of HMT, RMT, and three baseline models (OPT 2.7B, RWKV 3B, OpenLlamaV2 3B), evaluated over the PG-19 dataset. HMT outperforms RMT by 3.98% for OPT and 6.85% for OpenLlamaV2. For RWKV, HMT can boost the effectiveness by 9.96%.
6 Results and Key Observations
6.1 Impacts on Context-constrained Models
By applying an additional 0.5% $\sim$ 2% of parameters, HMT can augment any context-constrained models to process long-context inputs. We demonstrate this feature with general language modeling tasks and question-answering tasks.
HMT significantly improves the backbone models in general language modeling tasks when processing long inputs. Figure 4 and 5 compare the effectiveness of OPT 2.7B and OpenLlamaV2 3B models with and without HMT on the Wikitext-103 and PG-19 datasets. Over input spanning from 2k $\sim$ 100k tokens, HMT improves effectiveness by 25.5% for OPT and 17.6% for OpenLlamaV2 over the Wikitext-103 dataset. Moreover, for the PG-19 dataset, HMT enhances the effectiveness by 11.4% for OPT and 9.48% for OpenLlamaV2. In sum, HMT offers consistent benefits for context-constrained models.
HMT enhances long-answer contextual reasoning and short-answer prediction ability in question-answering tasks. One of the use cases of HMT is handling question-answering tasks that involve multiple contexts. Thus, we select the PubMedQA dataset and derive a long context QA benchmark to evaluate the effectiveness of HMT. We assess HMT with Llama 2 7B as the backbone model, shown in Figure 6 and 7. Two metrics are employed: for long answers, we compute the PPL to evaluate the contextual reasoning of HMT; for short answers, we measure the response accuracy. For samples with $M$ from 2 to 10, HMT increases the effectiveness in PPL by 9.48% for long answers. For short answer tasks, HMT is 1.0% more accurate than the backbone model. In sum, HMT boosts both the correctness and reasoning ability of context-constrained models in long-context QA tasks.
Figure 6: Effectiveness of RMT and HMT applied on Llama-2 7B, evaluated over PubMedQA dataset with modification for long context processing. The answer in each sample is the long answer. HMT is 8.98% more effective than RMT in PPL.
Figure 7: Response accuracy of RMT and HMT applied on Llama-2 7B, evaluated over PubMedQA dataset with modification for long context processing. The answer in each sample is the short answer. HMT is 1.8% more accurate than RMT on average.
6.2 Comparison to Long Context Models
Combined with context-constrained models, HMT can be more effective than long-context models. For example, HMT with OpenLlamaV2 3B is 44.0% and 14.9% more effective on Wikitext-103 and PG-19 respectively. Moreover, applying HMT to long-context models can further improve their effectiveness. HMT can enhance the effectiveness of RWKV by 16.5% on Wikitext-103 and 9.96% on PG-19 as shown in Figure 4 and 5. Beyond recurrent models, HMT also has better effectiveness with reduced memory consumption, indicating the usefulness of memory augmentation. For example, the AMD MI210 GPU cannot handle inferencing 30k token inputs with the Yi-6B-200K model due to memory constraints. Applying a sliding window strategy with a 5.2K-token window (Yi-6B-SW-5.2K), the model consumes 44.8 GB VRAM. On the contrary, HMT + Yi-6B-200K requires only 33.9 GB VRAM to process 30k tokens with a small segment length (512 tokens), with 2% effectiveness improvement. Table 4 presents the effectiveness of long-range models on Wikitext-103 compared with several HMT-augmented models, including Mamba and Mistral models.
Table 4: Effectiveness of long context models and HMT with various backbone models. The input size is 30k tokens and the dataset is Wikitext-103. HMT not only outperforms long context models when applied to context-constrained models but also can further improve the effectiveness of these models.
Model Max Context Test PPL (Wikitext)
RWKV 3B $\infty$ 13.13
Mamba 370M $\infty$ 87.08
Yi-6B-200K 200K OOM
Yi-6B-SW-5.2K 200K 6.89
Mistral-7B 32K 5.47
HMT + OPT 350M $\infty$ 13.04
HMT + OpenLlamaV2 3B $\infty$ 7.04
HMT + RWKV 3B $\infty$ 10.94
HMT + Mamba 370M $\infty$ 16.71
HMT + Yi-6B-200K $\infty$ 6.75
HMT + Mistral-7B $\infty$ 5.12
Furthermore, compared with other memory-augmented models, HMT is not only easy to use but also more effective. The Memorizing transformer (Wu et al., 2022) is a memory-augmented long-context model. This model requires training from scratch per downstream task before inference, which is not user-friendly for general usage. We compared the Memorizing Transformer trained on the Wikitext-103 dataset with HMT + OPT 350M, which has a similar parameter count, shown in table 5. With only 700 training steps, HMT outperforms the Memorizing Transformer trained with 10k steps by 2x in perplexity, indicating that HMT is more efficient and effective to use. We also compared HMT with another memory-augmented model, LongMem (Wang et al., 2024), shown in Appendix B.
Table 5: Effectiveness of Memorizing Transformer and HMT + OPT 350M models over Wikitext-103 dataset with 30k-token samples.
Model Test PPL (Wikitext)
Memorizing Transformer 31.51
HMT + OPT 350M 13.67
6.3 Comparison to RMT
Our assessment indicates that HMT is generally better at both language modeling and question-answering tasks than RMT.. Figure 4 and 5 compare the effectiveness between HMT and RMT with the Wikitext-103 and PG-19 datasets. HMT improves the effectiveness by 13% for OPT and 10.8% for OpenLlamaV2 on the Wikitext-103 dataset compared with applying RMT. Moreover, in contrast to RMT on the PG-19 dataset, HMT performs 3.98% better for OPT and 6.85% better for OpenLlamaV2.
For recurrent sequence models, HMT can further increase the effectiveness of RWKV while RMT will degrade the performance for both datasets, as demonstrated in Figure 5. Since RWKV has already compressed past tokens and passed hidden states along the sequence, applying RMT to RWKV is simply re-weighting past information compressed in states periodically. This was originally done by the time-mixing module of RWKV. Therefore, the advantage of memory augmentation is limited. Due to the gradient vanishing issue mentioned in Appendix H, the model is harder to train with RMT, leading to inferior performance. However, we believe that the memory recall mechanism in HMT helps RWKV to select previous hidden states with the most relevance, boosting the effectiveness.
Evaluating with the PubMedQA dataset, HMT improves the effectiveness in PPL by 8.98% for long answers compared with RMT. For short answer tasks, HMT is 1.8% more accurate than RMT on average.
6.4 Ablation Study
@@
We conduct ablation studies regarding the memory recall mechanism to demonstrate (1) memory recall is beneficial, (2) partial summarization of a segment in memory recall can speed up inference while maintaining similar effectiveness, and (3) caching more memory embedding can improve the effectiveness of HMT.
@@
Impact of memory recall mechanism. Figure 8 displays the advantages of having a memory recall mechanism in HMT for long context input with context switching. For any tested input length, the effectiveness of HMT with memory recall outperforms that without memory recall. Furthermore, when the memory recall mechanism is deployed, the effectiveness improves for the OPT 350M backbone model or tends to improve for the OPT 2.7B backbone model as the input sequence length grows, demonstrating better scalability of HMT.
@@
Figure 8: Effectiveness of HMT with and without memory recall mechanism for OPT 350M and 2.7B as the backbone models. The inputs are extracted from the Wikitext-103 dataset with up to 100k tokens.
Impact of summarizing partial segment in memory recall. To overlap or reduce the inference time of the previous segment with the representation extraction of the next segment, it is necessary to prefetch only the first $l$ tokens in the segment for summarization as described in Appendix G. In the experiment, we select half of the segment for representation extraction. We examine the model that extracts the whole segment and compare the effectiveness, depicted by Figure 9. The impact is negligible. We hypothesize that the start of a segment contains enough information about the overall topic for memory recall, which is intuitive as humans can capture the concepts by browsing keywords or segments instead of reading the whole paragraphs.
@@
Impact of limited cached memory embeddings. Due to memory constraints, we only cache the most recent 300 memory embeddings for memory recall. Figure 10 depicts the relationship between the number of cached memory embeddings and the effectiveness of HMT with Llama 2 7B over the Wikitext-103 dataset with 100k-token samples. We observed that increasing the window of cached memory benefits the effectiveness, but the improvement becomes marginal. We hypothesize that HMT is more likely to recall recent memory embeddings in the Wikitext-103 dataset. Figure 11 plots the frequency distribution of the distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory recall mechanism. 6.5% of the segments retrieve memory tokens within 2 segments. This indicates the importance of local context. However, the long context memory recall still exists. A possible explanation is that entries in Wikipedia may refer to other entries through hyperlinks and related context, and HMT discovers this long-context relationship and recalls the relevant information. For the PubMedQA dataset with multiple contexts concatenated, the long context recall also exists. We examine HMT with one sample containing 4 contexts illustrated in Figure 12. We inspect the highest softmax score of the memory recall cross attention when processing the segment that contains the questions and discover that it corresponds to the first context, which is the original context used for the question.
@@
Figure 11: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory recall mechanism. The dataset evaluated is the Wikitext-103.
Figure 12: Example of HMT recalling the correct context in PubMedQA dataset with 4 contexts.
7 Conclusion and Ongoing Works
We present HMT, a framework to augment LLMs’ long-range language processing ability with context switching. Inspired by the brain’s memory hierarchy, HMT imitates human memorization behavior by deploying hierarchical memorization and the memory recall mechanism. Compared with other models that can process infinitely long inputs, HMT has a more stable and better effectiveness for input with context switching. HMT can be directly applied to any pretrained fixed-context models with higher BPTT unroll depth than RMT. Moreover, it can boost other long-range transformer models which are already effective in processing infinitely long input. Our model not only provides LLM accessibility to resource-constrained applications but also represents a step forward to lifelong language processing tasks.
Currently, HMT will remove the oldest memory token when the total number of memory tokens stored is above a context window $N$. Hence, the model cannot recall more than $N$ segments. To recall any past segment, a possible solution is applying an extra hierarchy over the memory recall mechanism: we partition the memory tokens into segments and encode each segment into a tag vector. Memory recall requires encoding the current segment into a query tag and searching for the best memory token segment before computing the cross attention. Memory token segments can be stored throughout the machine (SSD, DRAM, and SRAM).
8 Ethical and Societal Impact
This work aims to advance the field of natural language processing. Specifically for HMT, it provides a foundation for building a lifelong AI assistant that can keep adapting to users’ habits by continuously monitoring users’ behavior, communicating with users, and memorizing experiences. Such an assistant offers convenience to people’s daily lives, while also raising concerns about privacy leakage. Nevertheless, with further efforts to deploy it on edge devices without network connections, this issue can be resolved.
Acknowledgement
Appendix A Other Related Works
The memory-augmented long-context transformer has been an active research topic in recent years. LongMem (Wang et al., 2024) chunks the long-document input into segments and caches the attention keys and values for each segment. During the inference of a segment, LongMem will select relevant key-value embedding pairs by computing the attention score between token embeddings and the cached key embeddings and fuse the top k embeddings. Memorizing Transformer (Wu et al., 2022) also caches the key-value embedding pairs similar to LongMem, but utilizes a kNN search to retrieve information similar to Unlimiformer. RET-LLM (Modarressi et al., 2023) employs prompt engineering to store the informative context in a database and search keywords when the context involves questions. While these works can precisely retrieve contexts, they are not scalable due to the increasing memory consumption of storing long contexts without compression. Segment-level recurrent models, such as EMMA (Moro et al., 2023), Memformer (Wu et al., 2020), and Transformer-XL (Dai et al., 2019), attempt to compress memory throughout the recurrence to reduce memory consumption. EMMA composes long-term memory from multiple short-term memory by linear combination and concatenates long and short-term memory to augment the segments. Transformer-XL propagates the compressed memory states derived from the attention of current layers to the previous layers for every iteration. Memformer augments the attention with the stored memory embeddings per time step and retrieves information using the cross-attention layer of the encoder-decoder model. However, Memformer employs a forgetting network to remove irrelevant context similar to LSTM, which can potentially delete useful contexts for unseen inputs. On the other hand, HMT compresses contexts into embeddings and retrieves information precisely without requiring a forgetting network to remove information permanently. Another issue with some of these works, including Memorizing Transformer, Memformer, Transformer-XL, and Unlimiformer, is that they need to fundamentally change the model architecture or inject new adapters based on different base model architecture. It makes deployment and extension to future LLMs very expensive. HMT avoids this issue by having a model-independent plug-and-play framework.
Appendix B Comparison to LongMem
Unlike HMT, LongMem (Wang et al., 2024) operates on the key and value caches of each layer of the model and requires caching long caches to capture distant context. To compare with LongMem, we pick the Qwen1.5-0.5B (Bai et al., 2023) model as the backbone model and train HMT by 700 steps with 4 segments over the ArXiv subset of the Pile dataset (Gao et al., 2020), which is the same setting as LongMem. Due to the large storage consumption of the training subset of the Pile dataset, we only extract the ArXiv subset in the validation and test split. HMT is trained on the validation set and tested on the test set. Table 6 illustrates that HMT + Qwen1.5-0.5B achieves lower PPL, with a smaller parameter size and in-memory length (number of memory embeddings, which is the number of key and value embeddings cached for LongMem). This indicates that HMT is memory efficient.
Table 6: Effectiveness of LongMem and HMT + Qwen1.5-0.5B models over ArXiv subset of the Pile dataset. With HMT, the Qwen1.5-0.5B model can obtain better effectiveness with fewer parameters and shorter memory, after 700 steps of update. The result for LongMem comes from the original paper.
Model # Params Segment Length In-memory Length Test PPL (ArXiv)
LongMem 558M 1k 65k 10.08
HMT + Qwen1.5-0.5B 463M 1k 300 9.06
Appendix C Comparison to Unlimiformer
There are two major differences between a previous retrieval-augmented model, Unlimiformer (Bertsch et al., 2023), and HMT in terms of the memory recall mechanism:
  • Unlimiformer retrieves the information with kNN search over the collection of encoded token segments, while HMT uses cross-attention. We believe there are several advantages of using cross-attention: (1) Attending Top K’s most similar token segments still introduces information loss. Regarding the self-attention layer, the aggregation of tokens with less similar encodings may positively contribute to the quality of the final output. On the other hand, cross-attention fuses all cached hidden embeddings, weighted by the relative similarity, which captures the whole context. (2) The output of the cross-attention is a single embedding, which has lower computational overhead compared to attending k extra tokens.
In terms of usage, Unlimiformer targets encoder-decoder models and injects retrieval modules into the backbone model. Although the authors recently added support for decoder-only models for token generation, only the Llama model architectures (Touvron et al., 2023) can be applied and the training/evaluation procedure is not specified. This is one of the biggest challenges for Unlimiformer to adapt to future LLMs for validation and generation. On the contrary, HMT focuses on decoder-only models. Since HMT does not inject new modules in the backbone model, it is effortless to adapt to future LLMs.
Appendix D Ablation Study of HMT Memory Recall Behavior

### v3

@@ Appendix D HMT, RMT, and Baseline Training Details and Hyperparameters
Table 7: Training and fine-tuning configurations for the backbone models (OPT 350M, Mamba 370M, OPT 2.7B, RWKV 3B, OpenLlamaV2 3B, Llama 2 7B, Yi-6B-200K, and Mistral 7B) and the modified model after applying RMT and HMT. S1 and S2 denote the first stage and the second stage of multi-stage training for HMT. 4 AMD MI210 GPUs cannot train larger models. The given learning rate is the starting learning rate and will decay by a factor of 0.9 for OPT, OpenLlamaV2, and RWKV models and 0.7 for the remaining models for every 100 steps. The batch size is 2.
@@
Model Input Length Segment Length Learning Rate Training Steps Extra Param
OPT 350M / SmolLM 135M 2048 1024 1e-5 700
@@
The baselines are assessed using the sliding window attention for context-constrained models to control the memory consumption for a fair comparison. Due to the limited VRAM of GPUs, we shrink the segment length for the larger backbone model. Also, increasing memory token size does not improve the effectiveness as Bulatov et al. (2022) suggested. Thus, both RMT and HMT apply memory tokens with a length of 1. For RMT, we utilize the maximum BPTT unroll depth with the best effectiveness that the GPUs can handle. HMT is trained with the multi-stage training technique illustrated in Section F. The first stage (S1) is trained with 2 segments, and the second stage (S2) is trained with the maximum BPTT unroll depth that the GPUs can manage. The size of long-term memory is 300 (i.e., $N=300$) for OPT, SmolLM, OpenLlama, Yi, and Mistral models and 400 ($N=400$) for the rest of the models to capture sufficient contexts and we summarize half of the segment for representation extraction (i.e., $j=L/2$). We observed that the benefit of increasing $N$ is diminishing and stops at 300 for 100k token inputs, described in Appendix E. We select the learning rate for HMT, RMT, and the baseline to optimize the effectiveness. Furthermore, HMT preserves 32 tokens ($k=32$) from the previous segment as the sensory memory. All models are trained with 700 steps, which is sufficient to converge the training loss. For models with HMT, we first pretrain the model with RedPajamaV2 Computer (2023) with 700 steps, then finetune the model for the downstream tasks with 700 steps. For LongBench, the evaluation metrics are the same as the original work and we use the Huggingface evaluate package to compute Rouge-L and F1 score. For parameter-efficient training, LoRA Hu et al. (2021) with rank 8 is applied to models with high training memory consumption (Llama 2 7B, Mamba 370M, Yi-6B-200K, and Mistral 7B). For the rest of the models, we finetune all parameters in the backbone model. All experiments are done with 3 random seeds and we take the average metrics.
Appendix E Ablation Study
@@
We conduct ablation studies regarding the memory retrieval mechanism to demonstrate that (1) memory retrieval is beneficial, (2) partial summarization of a segment in memory retrieval can speed up inference while maintaining similar effectiveness, and (3) caching more memory embedding can raise the effectiveness of HMT.
@@
Impact of memory retrieval mechanism. Figure 9 displays the advantages of having a memory retrieval mechanism in HMT for long context input with context switching. For any tested input length, the effectiveness of HMT with memory retrieval outperforms that without memory retrieval. Furthermore, when the memory retrieval mechanism is deployed, the effectiveness improves for the OPT 350M backbone model or tends to improve for the OPT 2.7B backbone model as the input sequence length grows, demonstrating better scalability of HMT.
@@
Figure 8: Effectiveness of HMT with and without the memory retrieval mechanism for OPT 350M and 2.7B as the backbone models. The inputs are extracted from the Wikitext-103 dataset with up to 100k tokens.
@@
Impact of summarizing partial segment in memory retrieval. To overlap or reduce the inference time of the previous segment with the representation extraction of the next segment, it is necessary to prefetch only the first $l$ tokens in the segment for summarization. In the experiment, we select half of the segment for representation extraction. We examine the model that extracts the whole segment and compare the effectiveness, depicted by Figure 9. The impact is negligible. We hypothesize that the start of a segment contains enough information about the overall topic for memory retrieval, which is intuitive as humans can capture the concepts by browsing keywords or segments instead of reading the whole paragraphs.
Impact of limited cached memory embeddings. Due to memory constraints, we only cache the most recent 300 memory embeddings for memory retrieval. Figure 11 depicts the relationship between the number of cached memory embeddings and the effectiveness of HMT with Llama 2 7B over the Wikitext-103 dataset with 100k-token samples. We observed that increasing the window of cached memory benefits the effectiveness, but the improvement becomes marginal. We hypothesize that HMT is more likely to recall recent memory embeddings in the Wikitext-103 dataset. Figure 11 plots the frequency distribution of the distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory retrieval mechanism. 6.5% of the segments retrieve memory tokens within 2 segments. This signals the importance of local context. However, the long-context memory retrieval still exists. A possible explanation is that entries in Wikipedia may refer to other entries through hyperlinks and related context, and HMT discovers this long-context relationship and recalls the relevant information.
In our experiments, we store the most recent 300 memory embeddings to balance the trade-off between retrieval effectiveness and computational efficiency. Theoretically, storing $\lceil M/L\rceil$ embeddings is sufficient to handle up to $M$ token inputs, ensuring robust performance. For longer inputs, high-quality processing is still achievable if recent contextual information is more relevant to the prompt.
@@
Figure 11: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory retrieval mechanism. The dataset evaluated is the Wikitext-103.
Figure 12: Training HMT + OPT 2.7B with the memory retrieval mechanism in two steps results in a better performance than using the mechanism to train HMT directly. Total training time is 902 s for multi-stage training and 1680 s for single-stage training on 4 AMD MI210 GPUs.
Appendix F Multi-stage Training
Since HMT introduces parameters for memory retrieval, we need to train new parameters and fine-tune the parameters of the backbone model cooperatively. Training HMT involves multiple segments of tokens to learn how to encode input tokens and retrieve information properly. Therefore, we split training HMT into two stages. In the first stage, The model is trained without the memory retrieval mechanism employing BPTT with 2 segments unrolled. BPTT saves the model checkpoint locally. Then, the memory retrieval mechanism loads and extends the train model in the second stage. At this point, BPTT trains the modified model by unrolling the maximum number of segments that the GPUs can handle to maximize the effectiveness, which is 15 in our experiment. Since the architecture of HMT is complex, breaking up training into two stages is beneficial for local optimization and improves long context inference performance compared with single-stage training. Figure 12 exhibits the performance difference between the multi-stage and single-stage training of the OPT 2.7B model with HMT for long-context inputs. Since Stage 1 involves a shorter training sequence length and a simpler recurrent architecture than Stage 2, training with Stage 1 is faster per iteration (1.15 s/iteration) than Stage 2 (3.36 s/iteration). Within the same number of training steps, multi-stage training obtains better effectiveness and lower total training time than single-stage training.
Appendix G HMT Memory Retrieval Behavior